# Cost Functions in Economic Complexity

### v2

@@ Abstract
Economic complexity algorithms aim to uncover the hidden capabilities that drive economic systems. Here, we present a fundamental reinterpretation of two of these algorithms, the Economic Complexity Index (ECI) and the Economic Fitness and Complexity (EFC), by reformulating them as optimization problems that minimize specific cost functions. We show that ECI computation is equivalent to finding eigenvectors of the network’s transition matrix by minimizing the quadratic form associated with the network’s Laplacian, thus revealing its limitations in capturing the diversification of countries. For EFC, we derive a novel cost function that exploits the algorithm’s intrinsic logarithmic structure and clarifies the role of its regularization parameter. Additionally, we establish the uniqueness of its solution, providing theoretical foundations for its application. This optimization–based reformulation bridges economic complexity and established frameworks in spectral theory, network science, and optimization. The theoretical insights translate into practical computational advantages: we introduce a conservative, gradient-based update rule that substantially accelerates algorithmic convergence, with potential implications for a broader class of algorithms, including the Sinkhorn–Knopp method. Beyond advancing our theoretical understanding of economic complexity indicators, this work opens new pathways for algorithmic improvements and extends applicability to general network structures beyond traditional bipartite economic networks.
Keywords Complex Networks  $\cdot$ Economic Complexity Index  $\cdot$ Economic Fitness and Complexity  $\cdot$ Cost Functions
1 Introduction
@@
Economic complexity measures are designed to characterize the productive capabilities of countries and the structure of international trade [1, 2, 3]. A first attempt to formalize this idea was made by Hidalgo and Hausmann through the Economic Complexity Index (ECI) [4, 5], based on a linear spectral method applied to the country-product network [6, 7]. The algorithm was later shown to be mathematically equivalent to a spectral clustering method that partitions countries and products into groups. Subsequently, Tacchella et al. [1, 8] proposed a different approach with the Economic Fitness and Complexity (EFC) method, introducing a non-linear iterative framework to better capture the interplay between country diversification and product competitiveness [9, 10]. The concept of fitness is closely linked to the notion of nestedness in bipartite networks [11], and has been shown to be a reliable predictor of GDP growth [10]. Later refinements of EFC, such as the regularized Non-Homogeneous Economic Fitness and Complexity (NHEFC) method [12], aimed to ensure robust convergence and improve its applicability to empirical data.
@@
While the Economic Complexity approach has demonstrated strong empirical relevance in understanding economic growth and trade networks, its mathematical foundations remain an active area of research. Previous studies [6, 7] have explored the connections between economic complexity indices and spectral methods in network science. In particular, [13] recasts both ECI and EFC within a common spectral framework based on bipartite networks, highlighting their structural similarities despite substantial differences in algorithmic formulation. Indeed, the two approaches produce markedly different country rankings, underscoring the distinct informational content captured by each method [9]. In [14, 15], the authors investigate the relation between economic complexity indices and standard notions of centrality, showing that these algorithms can provide valuable insight into the spectral properties of the underlying graph. More recently, ECI and EFC have also been shown to be formally equivalent to network centrality measures [16]. Extending these algorithms to arbitrary network structures, beyond the canonical bipartite formulation, has significantly broadened their applicability and interpretability. Therefore, a deeper theoretical understanding of these approaches is crucial to expanding their potential applications, as already demonstrated in fields such as ecology [17, 18, 19].
@@
This study demonstrates that both ECI and EFC can be reformulated as optimization problems, each minimizing a distinct cost function. Specifically, we show that the ECI algorithm minimizes an energy function derived from the network’s Laplacian matrix. In contrast, we prove that the cost function underlying EFC is strictly convex, thus providing a rigorous foundation for its optimization-based formulation. We also establish convergence to the fixed point (corresponding to the minimum of the cost function) in the case of regular graphs, and prove local convergence for general undirected and unweighted graphs. In addition, we propose a modification to the iterative algorithm that directly follows the gradient descent of the potential, significantly accelerating numerical convergence and potentially applicable to the Sinkhorn–Knopp class of algorithms for matrix renormalization.
The results presented in this paper enhance the interpretability of economic complexity algorithms and pave the way for methodological improvements and new applications in network-based data analysis. The reformulation of these indices as cost-minimization problems contributes to a more rigorous theoretical understanding of economic complexity and opens new perspectives for their application beyond traditional bipartite economic networks.
2 Materials and Methods
2.1 Data
@@
2.2 Previous work
In our previous work [16, 24], we described two methods for quantifying economic complexity: the Economic Complexity Index (ECI) and the Economic Fitness and Complexity (EFC). Both approaches are based on the country–product binary matrix $\mathbf{M}$. In this matrix, the entry $M_{cp}$ is set to 1 if country $c$ significantly exports product $p$, and 0 otherwise, as determined by the Revealed Comparative Advantage (RCA) criterion [31].
@@ 2.2 Previous work
After two iterations, these lead to the transition matrices $\mathbf{N}_{1}=\mathbf{D}^{-1}\mathbf{M}\mathbf{U}^{-1}\mathbf{M}^{T}$ and $\mathbf{N}_{2}=\mathbf{U}^{-1}\mathbf{M}^{T}\mathbf{D}^{-1}\mathbf{M}$ whose second eigenvectors (beyond the trivial constant principal eigenvector) capture the nontrivial complexity rankings [6, 7].
In contrast, EFC employs a non-linear map [1]. To ensure robust convergence, we used the regularized Non-Homogeneous version of EFC (NHEFC) given by [12]
$\begin{cases}{{F}}^{(n)}_{c}=\delta+\sum_{p=1}^{N_{p}}M_{cp}/{{P}}^{(n-1)}_{p}\\ {{P}}^{(n)}_{p}=\delta+\sum_{c=1}^{N_{c}}M_{cp}/{{F}}^{(n-1)}_{c}\end{cases}$ (1)
where $F^{(0)}_{c}=P^{(0)}_{p}=1$ and $\delta=10^{-3}$.
@@
thereby generalizing the approach, originally designed for bipartite structures, to any simple-graph structure. In Eq. (3), the notation $(\vec{V})^{-1}$ stands for taking the reciprocal of each element of the vector $\vec{V}$. This mono-partite formulation is a key generalization, allowing the ECI and EFC concepts to be applied to any arbitrary network structure, not just the bipartite country-product networks for which they were originally designed.
2.3 Cost Functions
@@ 2.3.1 ECI
$U(\vec{V})=\frac{1}{4}\sum_{i,j}A_{ij}(V_{i}-V_{j})^{2}=\dfrac{1}{2}\sum_{i,j}L_{ij}V_{i}V_{j}.$ (6)
The minimum of this function occurs when $V_{i}=V_{j}$ for all $i$ and $j$, which corresponds to the constant vector. This is also the first eigenvector of the Laplacian matrix, representing the solution that the ECI algorithm approaches after many iterations.
2.3.2 EFC
@@ 3.1 Convergence properties of ECI
Since this fixed point (the constant vector) does not provide a meaningful solution of the problem of ranking countries or products, the original formulation of ECI [4] prescribes stopping the iteration after a fixed number of steps. As discussed in several works [13, 24, 2], the intermediate steps of the iteration retain information about the subdominant eigenvectors, particularly the second one. As a result, ECI can be obtained in two equivalent ways: (i) by iterating the map Eq. (2) for a fixed number of steps, or (ii) by directly computing the second eigenvector of the transition matrix $\mathbf{K}^{-1}\mathbf{A}$. This eigenvector maximizes the partitioning of the network, in line with the interpretation of ECI as a spectral algorithm for community detection [13]. This observation further supports the evidence that ECI has intrinsic limitations in capturing the true diversification structure of countries [9].
3.2 Convergence properties of EFC
3.2.1 Uniqueness of solution via convexity analysis
@@
Here, $\mathbf{Q}$ is the signless Laplacian matrix of the weighted graph, defined by $Q_{ij}=\delta_{ij}S_{i}+W_{ij}$. From a well-established property in spectral graph theory, the signless Laplacian $\mathbf{Q}$ of any graph (weighted or unweighted) is positive semi-definite; that is, all of its eigenvalues are non-negative ($\lambda(\mathbf{Q})\geq 0$). The second term, $\mathbf{\Delta}$, is a diagonal matrix with entries $\Delta_{ii}=\delta e^{Z_{i}}$ and zeros elsewhere ($\Delta_{ij}=\delta_{ij}\delta e^{Z_{i}}$).
@@
where $N$ is the number of nodes and $E$ is the number of links in the network. This result demonstrates that $\delta$ effectively regulates the potential function, preventing divergence and ensuring that $U$ remains bounded.
@@
3.2.2 Solution in the case of two connected nodes
@@
3.2.3 Solution in the case of a regular graph
@@
The specific shape of the cost function for the EFC algorithm in the case of a regular graph is shown in Fig. 1a. The potential curves highlight the fixed point, corresponding to the function’s global minimum. These curves also clearly demonstrate the strict convexity of the cost function across the entire fitness space. Additionally, Fig. 2a displays the potential landscape in the two-dimensional phase space defined by the first two components of the fitness vector, $V_{1}$ and $V_{2}$, for a regular graph with $N=10$ and degree $K=9$. The figure shows the trajectories of the algorithm starting from various initial conditions, all of which converge to the fixed point, located at $V\approx\sqrt{K}=3$.
3.2.4 Local and general convergence for a general graph
@@
While a general proof of convergence is difficult for arbitrary network structures, this property can be investigated empirically by analyzing the behavior of multiple trajectories initialized from different starting points. We carried out this analysis for the Zachary Karate Club network, as shown in Fig. 2b. The results indicate that, regardless of the initial condition in the phase space, all trajectories converge to the same fixed point. Although a formal proof is lacking, this evidence suggests that global convergence may hold more generally.
3.2.5 Reformulation of the map via gradient descent
@@
4 Discussion
@@
The optimization framework also offers significant practical and methodological advancements. For EFC, we have shown that the dynamics expressed in logarithmic variables are irrotational. This implies that the most efficient computational path to the solution is a direct gradient descent of the potential function. This leads to a new “conservative map” (Eq. (12) in the original text) that converges significantly faster and avoids the period-2 oscillations characteristic of the original map. Moreover, given the formal connection between the EFC algorithm and the Sinkhorn–Knopp algorithm, used for matrix renormalization and the computation of doubly stochastic matrices, the conservative map introduced here may also provide an effective computational strategy within that broader class of problems. Future work should explore whether this gradient-based formulation can systematically improve the convergence speed, stability, and theoretical understanding of Sinkhorn–type algorithms, particularly in the context of optimal transport and large-scale network renormalization.
@@
Interpreting economic complexity algorithms through potential functions opens the way to a novel perspective on their application. In economic networks, the potential function provides an interpretable energetic landscape for each node, capturing both local and global structural forces. This framing of the cost function as a potential opens up a rich conceptual framework. Given the demonstrated strict convexity of the potential, the energetic landscape is characterized by a single, unique global minimum, representing the system’s sole optimal configuration. At any other point, the potential $U(\vec{V})$ quantifies the “stress” or deviation from this ideal state. This picture re-frames economic resilience not as the depth of a potential well, but as the steepness of the landscape around the minimum; a steeper potential implies stronger restorative forces that pull the system back to its optimum after a shock. Consequently, structural transformation is reconceptualized not as a jump between different stable states but as a dynamic change of the landscape itself, where an alteration in the network’s topology reshapes the cost function and shifts the location of the unique minimum. Finally, the geometry of this landscape can inform measures of systemic risk, as a “flatter” potential in certain directions would indicate a system with weak restorative forces, vulnerable to large and inefficient deviations from its optimal configuration. This representation could provide powerful insights into the organization and resilience of the economic fabric, offering a new tool for empirical investigations in regional development, economic geography, and beyond.
@@
Author contributions:
V.D.P.S. conceptualized and led the study, including methodology, validation, and initial draft writing. A.B. contributed to methods and software development and created all visualizations. P.B. contributed to the formal analysis. All authors collaborated on the formal analysis and the final manuscript. All authors have read and agreed to the published version of the manuscript.
Funding:
This work was funded by the City of Vienna, Municipal Department 7, and the Federal Ministry of the Republic of Austria for Climate Action, Environment, Energy, Mobility, Innovation and Technology as part of the project GZ 2021-0.664.668.
Acknowledgments:

### v3

@@ Abstract
Economic complexity algorithms aim to uncover the hidden capabilities that drive economic systems. Here, we present a fundamental reinterpretation of two of these algorithms, the Economic Complexity Index (ECI) and the Economic Fitness and Complexity (EFC), by reformulating them as optimization problems that minimize specific cost functions. We show that ECI computation is equivalent to finding eigenvectors of the network’s transition matrix by minimizing the quadratic form associated with the network’s Laplacian. For EFC, we derive a novel cost function that exploits the algorithm’s intrinsic logarithmic structure and clarifies the role of the regularization parameter in its non-homogeneous version. Additionally, we establish the existence and uniqueness of its solution, providing theoretical foundations for its application. This optimization-based reformulation bridges economic complexity and established frameworks in spectral theory, network science, and optimization. The theoretical insights translate into practical computational advantages: we introduce a conservative, gradient-based update rule that substantially accelerates algorithmic convergence, with potential implications for a broader class of algorithms, including the Sinkhorn-Knopp method. Finally, we apply the energetic framework to a real-world trade network, demonstrating how link-wise energy provides a direct way to identify structurally relevant and vulnerable regions of the export matrix, thus complementing and enriching standard economic complexity analyses. Beyond advancing our theoretical understanding of economic complexity indicators, this work opens new pathways for algorithmic improvements and extends applicability to general network structures beyond traditional bipartite economic networks.
I Introduction
@@
Understanding the productive capabilities of countries and the structure of their trade networks is fundamental to explaining patterns of economic growth and development. Economic complexity measures are designed to characterize the productive capabilities of countries and the structure of international trade [1, 2, 3]. A first attempt to formalize this idea was made by Hidalgo and Hausmann through the Economic Complexity Index (ECI) [4, 5], based on a linear spectral method applied to the country-product network [6, 7]. The algorithm was later shown to be mathematically equivalent to a spectral clustering method that partitions countries and products into groups. Subsequently, Tacchella et al. [1, 8] proposed a different approach with the Economic Fitness and Complexity (EFC) method, introducing a non-linear iterative framework to capture better the interplay between country diversification and product competitiveness [9, 10]. The concept of fitness is closely linked to the notion of nestedness in bipartite networks [11], and is a reliable predictor of GDP growth [10]. Later refinements of EFC, such as the regularized Non-Homogeneous Economic Fitness and Complexity (NHEFC) method [12], were introduced specifically to resolve the convergence instabilities observed in the original algorithm [13, 14]. The regularization parameter $\delta$ is crucial in this regard, as it ensures robust convergence and improves applicability to empirical data. Henceforth, we will consistently refer to the non-homogeneous version of the EFC algorithm, unless otherwise specified.
@@
While the Economic Complexity approach has demonstrated strong empirical relevance in understanding economic growth and trade networks, its mathematical foundations remain an active area of research. Several studies have explored connections between economic complexity indices and established network science methods [6, 7]. Sciarra et al. [15] recast both ECI and EFC within a common spectral framework, highlighting their structural similarities despite algorithmic differences that produce markedly different country rankings [9]. Other work has investigated the relationships between these indices and standard centrality measures [16, 17, 18], demonstrating how economic complexity algorithms provide insight into the spectral properties of underlying graphs. Extending these algorithms to arbitrary network structures, beyond the canonical bipartite formulation, has significantly broadened their applicability and interpretability. Therefore, a deeper theoretical understanding of these approaches is crucial to expanding their potential applications, as already demonstrated in fields such as ecology [19, 20, 21].
@@
This study provides a rigorous optimization-theoretic foundation for both ECI and EFC algorithms. We demonstrate that each minimizes a distinct cost function: ECI is associated with an energy function derived from the network’s Laplacian matrix, while EFC is associated with a strictly convex function, for which we provide explicit characterization. We establish global convergence to the fixed point for regular graphs and prove local convergence with explicit rates for general undirected, unweighted graphs. Based on this analysis, we propose a gradient-descent modification that significantly accelerates numerical convergence, with potential applications to the broader class of Sinkhorn-Knopp matrix renormalization algorithms. Finally, we complement the theoretical analysis with an empirical study of the energetic landscape on a real-world trade network, showing how the two cost functions distribute energy across links and revealing structurally critical regions of the network. This link-level perspective provides both descriptive insight and practical guidance for assessing the organization and fragility of empirical economic systems.
The results presented in this paper enhance the interpretability of economic complexity algorithms, paving the way for methodological improvements and new applications in network-based data analysis. The reformulation of these indices as cost-minimization problems contributes to a more rigorous theoretical understanding of economic complexity and opens new perspectives for their application beyond traditional bipartite economic networks.
II Materials and Methods
II.1 Data
@@
For the empirical application, we use international trade data from the United Nations COMTRADE database  [^1] https://comtradeplus.un.org, classified according to the Standard International Trade Classification (SITC) Revision 2 [34]. We extract export flows for the year 2015 and construct the country–product bipartite matrix using the standard RCA threshold [35]. SITC Rev. 2 aggregates products into roughly 200 categories, providing a stable and widely used representation of countries’ export. This dataset offers a heterogeneous and realistic benchmark for assessing the energetic landscape of ECI and Fitness on real-world bipartite networks.
II.2 Previous work
In our previous work [18, 26], we described two methods for quantifying economic complexity: the Economic Complexity Index (ECI) and the Economic Fitness and Complexity (EFC). Both approaches traditionally operate on the country–product binary matrix $\mathbf{M}$. In this matrix, the entry $M_{cp}$ is set to 1 if country $c$ significantly exports product $p$, and 0 otherwise, as determined by the Revealed Comparative Advantage (RCA) criterion [35].
@@ II.2 Previous work
After two iterations, these expressions lead to the transition matrices $\mathbf{N}_{1}=\mathbf{D}^{-1}\mathbf{M}\mathbf{U}^{-1}\mathbf{M}^{T}$ and $\mathbf{N}_{2}=\mathbf{U}^{-1}\mathbf{M}^{T}\mathbf{D}^{-1}\mathbf{M}$ whose second eigenvectors (beyond the trivial constant principal eigenvector) capture the nontrivial complexity rankings [6, 7].
In contrast, EFC employs a non‐linear map [1]. To ensure robust convergence, we used the regularized Non-Homogeneous version of EFC (NHEFC) given by [12]
$\cases{{F}}^{(n)}_{c}=\delta+\sum_{p=1}^{N_{p}}M_{cp}/{{S}}^{(n-1)}_{p}\\ {{S}}^{(n)}_{p}=\delta+\sum_{c=1}^{N_{c}}M_{cp}/{{F}}^{(n-1)}_{c}$ (1)
where $F^{(0)}_{c}=S^{(0)}_{p}=1$ and $\delta=10^{-3}$.
@@
thereby generalizing the approach, originally designed for bipartite structures, to any simple-graph structure. In Eq. (3), the notation $(\vec{V})^{-1}$ stands for taking the reciprocal of each element of the vector $\vec{V}$. This mono-partite formulation is a key generalization, allowing the ECI and EFC concepts to be applied to any arbitrary network structure, not just the bipartite country-product networks for which they were originally designed [26].
II.3 Cost Functions
@@ II.3.1 ECI
[…]
The minimum of this function occurs when $V_{i}=V_{j}$ for all $i$ and $j$, which corresponds to the constant vector. This is also the first eigenvector of the Laplacian matrix, representing the solution that the ECI algorithm approaches after many iterations. This cost function has a natural physical interpretation: it corresponds to the potential energy of a networked system of one-dimensional springs, where nodes represent masses and edges represent springs connecting them.
II.3.2 EFC
@@ III.1 Convergence properties of ECI
Since this fixed point (the constant vector) does not provide a meaningful solution of the problem of ranking countries or products, the original formulation of ECI [4] prescribes stopping the iteration after a fixed number of steps. As discussed in several works [15, 26, 2], the intermediate steps of the iteration retain information about the subdominant eigenvectors, particularly the second one. As a result, ECI can be obtained in two equivalent ways: (i) by iterating the map Eq. (2) for a fixed number of steps, or (ii) by directly computing the second eigenvector of the transition matrix $\mathbf{K}^{-1}\mathbf{A}$. This eigenvector maximizes the partitioning of the network, in line with the interpretation of ECI as a spectral algorithm for community detection [15]. This observation clarifies the geometric nature of the algorithm: ECI minimizes the cut of the network, tending to group countries with similar export baskets together. This clustering behavior explains why ECI is mathematically distinct from, and orthogonal to, measures intended to capture diversification capabilities [15].
III.2 Convergence properties of EFC
III.2.1 Uniqueness of solution via convexity analysis
@@
Here, $\mathbf{Q}$ is the signless Laplacian matrix of the weighted graph, defined by $Q_{ij}=\delta_{ij}S_{i}+W_{ij}$. From a well-established property in spectral graph theory, the signless Laplacian $\mathbf{Q}$ of any simple graph (weighted or unweighted) is positive semi-definite; that is, all of its eigenvalues are non-negative ($\lambda(\mathbf{Q})\geq 0$). The second term, $\mathbf{\Delta}$, is a diagonal matrix with entries $\Delta_{ii}=\delta e^{Z_{i}}$ and zeros elsewhere ($\Delta_{ij}=\delta_{ij}\delta e^{Z_{i}}$).
@@
where $N$ is the number of nodes and $E$ is the number of links in the network. This result demonstrates that $\delta$ effectively regulates the potential function, preventing divergence and ensuring that $U$ remains bounded. From an optimization perspective, this explains why the non-homogeneous map does not suffer from the pathological convergence issues known to affect the original fitness algorithm. The regularization ensures the potential is strictly convex, guaranteeing a unique stable solution even for non-bipartite or non-nested topologies where the original map would fail.
@@
III.2.2 Solution in the case of two connected nodes
@@
III.2.3 Solution in the case of a regular graph
@@
The specific shape of the cost function for the EFC algorithm in the case of a regular graph is shown in Fig. 1a. The potential curves highlight the fixed point, corresponding to the function’s global minimum. These curves also clearly confirm the strict convexity of the cost function across the entire fitness space. Additionally, Fig. 2a displays the potential landscape in the two-dimensional phase space defined by the first two components of the fitness vector, $V_{1}$ and $V_{2}$, for a regular graph with $N=10$ and degree $K=9$. The figure shows the trajectories of the algorithm starting from various initial conditions, all of which converge to the fixed point, located at $V\approx\sqrt{K}=3$.
III.2.4 Local and general convergence for a general graph
@@
While a general proof of convergence is difficult for arbitrary network structures, this property can be investigated empirically by analyzing the behavior of multiple trajectories initialized from different starting points. We carried out this analysis for the Zachary Karate Club network, as shown in Fig. 2b. The results indicate that, regardless of the initial condition in the phase space, all trajectories converge to the same fixed point. Although a formal proof is lacking, this evidence suggests that global convergence may hold more generally. At the same time, we cannot rule out the existence of non-stationary orbits (e.g., periodic ones) that attract the trajectories of Eq. (4) for suitable choices of the initial configuration.
III.2.5 Reformulation of the map via gradient descent
Figure 4: Energetic landscape of ECI and EFC on the trade network. Link-wise energy for the country–product matrix (UN COMTRADE, SITC Rev. 2, 2015), with rows and columns sorted by the corresponding scores. For visualization purposes, energy values are smoothly interpolated within the non-zero “frontier” of the matrix, i.e., the region to the left of the boundary beyond which all entries are zero. (a) ECI energy, $E_{cp}\propto M_{cp}(ECI_{c}-ECI_{p})^{2}$, is lowest along the diagonal, consistent with its Laplacian structure, and increases for links between nodes with distant spectral positions. (b) EFC energy, $E_{cp}\propto M_{cp}/(F_{c}S_{p})$, concentrates along the “frontier” of the matrix, highlighting structurally critical links involving low-fitness countries or low-simplicity products. ECI thus reflects smooth spectral variation, while EFC identifies points of fragility more relevant for targeted perturbations.
@@ III.2.5 Reformulation of the map via gradient descent
It is worth noting that, by introducing a scaling factor $\eta$ in front of the gradient in Eq. (12), the basin of attraction of the fixed point under this modified dynamics expands as $\eta$ decreases, eventually covering the entire phase space in the limit $\eta\to 0$. Consequently, the robustness of the algorithm can be enhanced arbitrarily.
@@
III.2.6 Application to real trade networks
[…]
Panel (a) shows that the ECI energy landscape reproduces the geometry of a Laplacian quadratic form: the lowest-energy links lie close to the diagonal, where countries and products have similar ECI values, while high-energy contributions appear far from it. This structure reflects the spectral nature of the ECI cost function, which penalizes connections between nodes occupying distant positions in the eigenvector-based embedding.
The EFC landscape in panel (b) exhibits a different and more heterogeneous structure. High-energy links concentrate along the “frontier” of the matrix, typically involving low-fitness countries or low-simplicity products. This is consistent with the nonlinear amplification mechanism of the EFC algorithm: rare connections that tie highly capable countries to complex products, or peripheral countries to their few essential exports, contribute disproportionately to the potential. These links are structurally critical, since their removal would rapidly isolate low-fitness countries or disconnect the most complex products from the network [18].
This energetic perspective naturally suggests attack or pruning strategies based on link energy. Removing the highest-energy links under the ECI potential would mainly affect off-diagonal connections, but these links are typically redundant for highly diversified countries or sparsely populated for low-ECI products, resulting in limited structural impact. In contrast, removing high-energy links according to the EFC potential would directly target the structural weak points of the network: the essential exports of low-fitness countries and the few connections sustaining highly complex products. Such a strategy would more rapidly destabilize the system, potentially eliminating entire nodes. Hence, the Fitness energy identifies connections whose removal is expected to produce the most pronounced effects on the resilience of the trade network.
IV Discussion
@@
The optimization framework also offers significant practical and methodological advancements. For EFC, we have shown that the dynamics expressed in logarithmic variables are irrotational. This implies that the most efficient computational path to the solution is a direct gradient descent of the potential function. This leads to a new “conservative map” (Eq. (12) in the original text) that converges significantly faster and avoids the sawtooth oscillations characteristic of the original map. Moreover, given the formal connection between the EFC algorithm and the Sinkhorn-Knopp algorithm, used for matrix renormalization and the computation of doubly stochastic matrices, the conservative map introduced here may also provide an effective computational strategy within that broader class of problems. Future work should explore whether this gradient-based formulation can systematically improve the convergence speed, stability, and theoretical understanding of Sinkhorn-type algorithms, particularly in the context of optimal transport and large-scale network renormalization.
Interpreting economic complexity algorithms through potential functions opens the way to a novel perspective on their application. In economic networks, the potential function provides an interpretable energetic landscape, capturing both local and global structural forces. As illustrated in Fig. 4, the analysis of a real-world trade network shows how the energetic framework translates directly into empirical insight. The distribution of link-wise energy highlights structurally critical regions of the export matrix, revealing which connections sustain the organization and resilience of the system. By identifying links that contribute significantly to the potential, this approach offers a quantitative tool for diagnosing fragility and understanding how local disruptions may propagate through the network.
@@
Viewed from a broader perspective, these empirical patterns also motivate a more general interpretation of the potential function as a device for assessing systemic stability and the network’s response to shocks. Given the demonstrated strict convexity of the potential, the energetic landscape is characterized by a single, unique global minimum, representing the system’s sole optimal configuration. At any other point, the potential $U(\vec{V})$ quantifies the “stress” or deviation from this ideal state. This picture re-frames economic resilience not as the depth of a potential well, but as the steepness of the landscape around the minimum; a steeper potential implies stronger restorative forces that pull the system back to its optimum after a shock. Consequently, structural transformation is reconceptualized not as a jump between different stable states but as a dynamic change of the landscape itself, where an alteration in the network’s topology reshapes the cost function and shifts the location of the unique minimum. Finally, the geometry of this landscape can inform measures of systemic risk, as a “flatter” potential in certain directions would indicate a system with weak restorative forces, vulnerable to large and inefficient deviations from its optimal configuration. This representation could provide powerful insights into the organization and resilience of the economic fabric, offering a new tool for empirical investigations in regional development, economic geography, and beyond.
@@
Author contributions: V.D.P.S. conceptualized and led the study, including methodology, validation, and initial draft writing. A.B. contributed to methods and software development and created all visualizations. P.B. contributed to the formal analysis. All authors collaborated on the formal analysis and the final manuscript. All authors have read and agreed to the published version of the manuscript.
Funding: This work was co-financed by the Austrian Federal Ministry for Innovation, Mobility and Infrastructure (BMIMI) as part of the project GZ 2021-0.664.668 and the Oesterreichische Nationalbank as part of their funding programme for Austrian economic research.
Conflicts of interest: The authors declare no conflict of interest.
Abbreviations: The following abbreviations are used in this manuscript:
Acknowledgements.